O 8/19 points Answers identical conducting spheres originally have charges of -2.23 nc and 6.93 nc. If they were to touch and then separate again, what would be the charge on each sphere? The charge on each sphere would be Your response differs from the correct Enter a number with unlts 1 00%.
Answer #1

total charge = 6.93-2.23 = 4.70 nc

since spheres are identical

so charge on each = 4.70/2 = 2.35 nc